OpenStudy (anonymous):

what is 2=log 64 8 what is the equivalent exponential form of the equations?

OpenStudy (anonymous):

the 8 is suppose to be after the log but lower idk how to type that

OpenStudy (anonymous):

8^2=64

OpenStudy (lgbasallote):

\(\large \log_{8} 64?\) type in your equations editor \log_{8} 64

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

^that was how to make it appear that was not an answer lol..jsyk

OpenStudy (anonymous):

do you know the answer

OpenStudy (lgbasallote):

like @optiquest said 8^2 = 64 \(\large \log_{8} 64 = \log_{8} 8^2\) guess you know some logarithm properties? if not check this out.. http://openstudy.com/study#/updates/4fa45f8fe4b029e9dc34e0b5

OpenStudy (lgbasallote):

and yes i do know the answer lol

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

looks 8 posts up

OpenStudy (anonymous):

ok so i just type in that whole thing 8^2=64